# Orthogonal Howe duality and dynamical (split) symmetric pairs

### v2

@@ Acknowledgments
This project started as a collaboration with Hunter Dinkins and we are thankful for his input. We benefited from discussions with Pavel Etingof, Valerio Toledano Laredo, Leonid Rybnikov, Jasper Stokman, and Milen Yakimov. Thanks also to Stefan Kolb, and Weiqiang Wang for helpful correspondence about the rank $1$ formulas discussed in Remark 3.9.
@@ 2.1. Cartan involution
Let $\mathfrak{g}_{\mathbb{R}}$ be a real simple Lie algebra with Killing form $\kappa_{\mathbb{R}}(-,-)$. There is a Cartan involution $\theta_{\mathbb{R}}:\mathfrak{g}_{\mathbb{R}}\rightarrow\mathfrak{g}_{\mathbb{R}}$, which by definition is a Lie algebra automorphism of order $2$, such that $-\kappa_{\mathbb{R}}(-,\theta_{\mathbb{R}}(-))$ is positive definite. Cartan involutions are unique up to conjugation by an inner automorphism of $\mathfrak{g}_{\mathbb{R}}$.
@@
where $\mathfrak{k}_{\mathbb{R}}=\mathfrak{g}_{\mathbb{R}}^{\theta}$ and $\theta_{\mathbb{R}}$ acts on $\mathfrak{p}_{\mathbb{R}}$ by $-1$.
We further assume that the Lie algebra $\mathfrak{g}_{\mathbb{R}}$ is split, i.e. $\mathfrak{p}_{\mathbb{R}}$ contains a Cartan subalgebra $\mathfrak{h}_{\mathbb{R}}\subset\mathfrak{g}_{\mathbb{R}}$. The restriction of $\kappa_{\mathbb{R}}(-,-)$ to $\mathfrak{h}_{\mathbb{R}}$ is positive definite. In particular, $\kappa_{\mathbb{R}}(-,-)$ identifies $\mathfrak{h}_{\mathbb{R}}$ with $\mathfrak{h}_{\mathbb{R}}^{*}$ and there is an induced inner product $\kappa^{*}(-,-)$ on $\mathfrak{h}_{\mathbb{R}}^{*}$. Write $\langle-,-\rangle_{\mathbb{R}}$ for the natural pairing between $\mathfrak{h}_{\mathbb{R}}^{*}$ and $\mathfrak{h}_{\mathbb{R}}$.
Let $\mathfrak{g},\mathfrak{k},\mathfrak{p}$, and $\mathfrak{h}$ denote the complexifications of $\mathfrak{g}_{\mathbb{R}}$, $\mathfrak{k}_{\mathbb{R}}$, $\mathfrak{p}_{\mathbb{R}}$, and $\mathfrak{h}_{\mathbb{R}}$. We have $\theta=\mathbb{C}\otimes\theta_{\mathbb{R}}$, an order two automorphism of $\mathfrak{g}$ such that $\mathfrak{g}^{\theta}=\mathfrak{k}$ and $\mathfrak{p}$ is the $-1$ eigenspace of $\theta$. We also write $\kappa(-,-)$ and $\langle$-,$\rangle$ - , - ⟩ for the complexifications of $\kappa_{\mathbb{R}}(-,-)$ and $\langle-,-\rangle_{\mathbb{R}}$.
@@ 2.2. Roots and generators
For $\alpha\in R$, we denote by $t_{\alpha}\in\mathfrak{h}$ the element corresponding to $\alpha\in\mathfrak{h}^{*}$ under the identification $\mathfrak{h}\xrightarrow{\cong}\mathfrak{h}^{*}$ induced by $\kappa$, i.e. $t_{\alpha}\in\mathfrak{h}$ is the unique element such that
@@
This means $\alpha$ is real valued on $\mathfrak{h}_{\mathbb{R}}^{*}$ and $t_{\alpha}\in\mathfrak{h}_{\mathbb{R}}$. We fix generators $E_{\alpha}\in\mathfrak{g}_{\mathbb{R}}\cap\mathfrak{g}_{\alpha}$, for all $\alpha\in R$, such that $[E_{\alpha},E_{-{\alpha}}]=t_{\alpha}$, i.e. $\kappa_{\mathbb{R}}(E_{\alpha},E_{-\alpha})=1$, and $\theta_{\mathbb{R}}(E_{\alpha})=-E_{-\alpha}$ (it is possible by [Humphreys, §25.2]). Note that
@@
$[E_{\alpha},F_{\alpha}]=t_{\alpha}.$
@@ 2.3. Iwasawa decomposition
For $\mathfrak{sl}_{2}$, we denote the generator of $\mathfrak{k}=\mathfrak{so}_{2}$ simply by $B:=F-E$.
@@ 2.4. Chevalley involution and generators
Thus, $\theta$ is conjugate to the standard Chevalley involution, which swaps $e_{\alpha_{i}}$ and $f_{\alpha_{i}}$ and acts as $-1$ on $\mathfrak{h}$, the existence of which is a Corollary of Serre’s presentation of $\mathfrak{g}$. In general, Chevalley involution refers to the automorphism of $\mathfrak{g}$ corresponding to the Cartan involution of the split real form.
@@ 2.5. Integrable representations
The Lie algebra $\mathfrak{k}$ is reductive which follows from the classification result [KnappLieGroups, Appendix C]. Fix a connected reductive group $K$ with the Lie algebra $\mathfrak{k}$. We call a $\mathfrak{k}$-representation integrable if it comes from a representation of $K$. By a “finite-dimensional representation of $\mathfrak{g}$” we understand a Harish-Chandra $(\mathfrak{g},K)$-module, i.e. a $\mathfrak{g}$-module which is integrable over $\mathfrak{k}$. In what follows, we will not refer to $K$ explicitly.
@@ 2.6. Weights and representations
This is a non-degenerate $\mathfrak{g}$-invariant form on $\mathfrak{g}$. We also write (,)(-,-)( - , - ) to denote the induced form on $\mathfrak{h}^{*}$, which is $W$-invariant and for any short root $\alpha$ satisfies
@@
In fact, since $W$ acts irreducibly on $\mathfrak{h}^{*}$ these two properties characterize (,)(-,-)( - , - ).
@@ 3.1. Braid group action for Weyl group
lying in a certain completion of $\mathrm{U}(\mathfrak{sl}_{2})$. Its action is well-defined on any finite-dimensional $\mathfrak{sl}_{2}$-module $V$. Moreover, the Weyl group operator is group-like, i.e. $\Delta(\tilde{s}^{\mathfrak{sl}_{2}})=\tilde{s}^{\mathfrak{sl}_{2}}\otimes\tilde{s}^{\mathfrak{sl}_{2}}$.
@@ 3.2. Braid group action for boundary Weyl group: $\mathfrak{so}_{2}\subset\mathfrak{sl}_{2}$
Consider the $\mathfrak{sl}_{2}$-module $V(1)$, with weight basis $v_{+}$ and $v_{-}$. We have $b=f-e\in\mathfrak{so}_{2}$ acting by $b\cdot v_{+}=v_{-}$ and $b\cdot v_{-}=-v_{+}$. Thus, $V(1)|_{\mathfrak{so}_{2}}=\mathbb{C}\{v_{+}-iv_{-}\}\oplus\mathbb{C}\{v_{+}+iv_{-}\}\cong\mathbb{C}_{i}\oplus\mathbb{C}_{-i}$. In this representation, we have $\tilde{s}=b$, so $\tilde{s}$ acts on $\mathbb{C}_{i}$ by $i$ and on $\mathbb{C}_{-i}$ by $-i$.
@@
For any finite-dimensional module $V$ of $\mathfrak{sl}_{2}$, the action of $\tilde{s}^{\mathfrak{sl}_{2}}$ is equal to the action of $\exp\left(\frac{\pi b}{2}\right)$ understood in the sense of 3.4.
@@
By Example 3.3, the actions agree on $V(1)$. Since $\Delta(b)=b\otimes 1+1\otimes b$ is a sum of two commuting operators, the operator $\exp\left(\frac{\pi b}{2}\right)$ is group-like, i.e. $\Delta(\exp\left(\frac{\pi b}{2}\right))=\exp\left(\Delta(\frac{\pi b}{2})\right)=\exp\left(\frac{\pi b}{2}\right)\otimes\exp\left(\frac{\pi b}{2}\right)$. Since the operator $\tilde{s}$ is group-like as well, their actions agree on any tensor product $V(1)^{\otimes n}$. As any irreducible representation of $\mathfrak{sl}_{2}$ can be realized as a subrepresentation of a suitable tensor product, the claim is then a consequence of complete reducibility of finite dimensional $\mathfrak{sl}_{2}$-modules. ∎
@@
Experts in the theory of $\iota$quantum groups will be familiar with the above formulas. Non-experts should compare the divided power formulas in Definition 3.8 with the $\iota$divided power formulas in [Bao-Wang-new-approach, Conjecture 4.13] (the conjecture that this is the $\iota$canonical basis is proven in [Berman-Wang-formulas-for-divided]) and compare the formulas for the operators in equation (3.3) with [Zhang-Weinan-thesis, Section 16.3].
@@
We see that $s_{0}(2x)$ and $-is_{1}(2x+1)$ satisfy the same inductive relations as $\tilde{s}^{\mathfrak{so}_{2}}(2x)$ and $\tilde{s}^{\mathfrak{so}_{2}}(2x+1)$. Therefore, the statement follows from the equalities: $\tilde{s}^{\mathfrak{so}_{2}}(0)=s_{0}(0)=1$ and $\tilde{s}^{\mathfrak{so}_{2}}(1)=-is_{1}(1)=-i$. ∎
@@ 3.3. Braid group action for boundary Weyl group: $\mathfrak{k}\subset\mathfrak{g}$
Similarly to the usual Weyl group operators, we define the boundary Weyl group for the split symmetric pair $(\mathfrak{g},\mathfrak{k})$ in terms of simple reflection. In what follows, we are in the setting of Section 5.
@@
Let $\alpha\in\mathfrak{h}^{*}$ be a simple root. Define the boundary Weyl group operator $\tilde{s}^{\mathfrak{k}}_{\alpha}:=\exp\left(\frac{\pi b_{\alpha}}{2}\right)$, understood in the sense of 3.4. When $\alpha=\alpha_{i}$, we will also denote $\tilde{s}^{\mathfrak{k}}_{i}:=\tilde{s}^{\mathfrak{k}}_{\alpha_{i}}$.
@@
For a finite dimensional $\mathfrak{g}$-module $V$, we have the usual Weyl group operators $\widetilde{w}^{\mathfrak{g}}$ acting on $V$. By Proposition 3.5, the action of $\tilde{s}^{\mathfrak{k}}_{i_{1}}\ldots\tilde{s}^{\mathfrak{k}}_{i_{l}}$ on $V|_{\mathfrak{k}}$ coincides with the action of $\widetilde{w}^{\mathfrak{g}}$. In particular, if $X$ is an integrable $\mathfrak{k}$-module and $V$ is a finite dimensional $\mathfrak{g}$-module, then, recalling the conventions for $\Delta$ set forth in Section 2.5, we find
@@
Since the operators $\tilde{s}_{\alpha}^{\mathfrak{g}}$ satisfy the braid relations, it is enough to prove the proposition for integrable representations $X_{1},\ldots,X_{l}$ of $\mathfrak{k}$ such that any other irreducible representation can be obtained as a direct summand of $X_{i}\otimes V$ for some finite dimensional $\mathfrak{g}$-representation $V$. Moreover, it is enough to prove this for rank 2 algebras. This is done in Appendix B. ∎
@@
Moreover, Proposition 3.5 implies $\mathrm{Ad}_{\tilde{s}^{\mathfrak{g}}_{i}}=\mathrm{Ad}_{\tilde{s}^{\mathfrak{k}}_{i}}$, hence $\mathrm{Ad}_{\tilde{s}^{\mathfrak{g}}_{i}}$ commutes with the Cartan involution $\theta:\mathfrak{g}\rightarrow\mathfrak{g}$. Therefore,
@@ 4.1. Fusion operator
In what follows, we will often need to restrict to a certain open subset of $\mathfrak{h}^{*}$.
@@
We denote by $\mathcal{M}(\mathfrak{h}^{*})$ the space of meromorphic functions on $\mathfrak{h}^{*}$ with poles on the hyperplanes $\{\langle\alpha^{\vee},\lambda\rangle=k\}$ for all $\alpha\in R$ and $k\in\mathbb{Z}$.
@@
Let $V_{1},V_{2},\dots V_{d}$ be $\mathfrak{g}$-representations. For a function $f:\mathfrak{h}^{*}\rightarrow\mathrm{End}(V_{1}\otimes V_{2}\otimes\cdots\otimes V_{d})$, denote by $\lambda\mapsto f(\lambda-h|_{V_{i}})$, the function
@@
Let $V$ be an finite-dimensional representation of $\mathfrak{g}$. For a weight $\nu\in\mathfrak{h}^{*}$, we denote by $V[\nu]$ the corresponding weight space of $V$. Consider the space of intertwiners $\mathrm{Hom}_{\mathfrak{g}}(M_{\lambda},M_{\mu}\otimes V)$.
@@
    $\Phi^{v}_{\lambda}(v)=\Xi(\lambda)(m_{\mu}\otimes v).$
@@
The second assertion follows directly from [EtingofStyrkas, Proposition 2.1]: for every $\mu$, the singular vectors in $M_{\mu}\otimes V$ satisfying (4.3) can be obtained by applying the inverse of the Shapovalov form. Since the matrix coefficients of the latter are rational functions in $\mu$ and $\mu=\lambda-h|_{V}$ in the notations of 4.4, we conclude. ∎
@@
defines an intertwiner. Assuming the weights $\lambda$, $\mu$, and $\nu$ are generic, this intertwiner can be identified with $\Phi_{\lambda}^{u}$ for some $u\in W\otimes V[\lambda-\nu]$.

### v3

@@ Acknowledgments
This project started as a collaboration with Hunter Dinkins and we are thankful for his input. We benefited from discussions with Pavel Etingof, Valerio Toledano Laredo, Leonid Rybnikov, Jasper Stokman, and Milen Yakimov. Thanks also to Stefan Kolb, and Weiqiang Wang for helpful correspondence about the rank $1$ formulas discussed in Remark 3.9. E.B. was partially supported by the NSF MSPRF-2202897.
@@ 2.1. Cartan involution
Let $\mathfrak{g}_{\mathbb{R}}$ be a real simple Lie algebra with Killing form $\kappa_{\mathbb{R}}(-,-)$. There is a Cartan involution $\theta_{\mathbb{R}}:\mathfrak{g}_{\mathbb{R}}\rightarrow\mathfrak{g}_{\mathbb{R}}$, which by definition is a Lie algebra automorphism of order $2$, such that $\kappa_{\mathbb{R}}(-,\theta_{\mathbb{R}}(-))$ is positive definite. Cartan involutions are unique up to conjugation by an inner automorphism of $\mathfrak{g}_{\mathbb{R}}$.
@@
where $\mathfrak{k}_{\mathbb{R}}=\mathfrak{g}_{\mathbb{R}}^{\theta}$ and $\theta_{\mathbb{R}}$ acts on $\mathfrak{p}_{\mathbb{R}}$ by 1-1.
We further assume that the Lie algebra $\mathfrak{g}_{\mathbb{R}}$ is split, i.e., $\mathfrak{p}_{\mathbb{R}}$ contains a Cartan subalgebra $\mathfrak{h}_{\mathbb{R}}\subset\mathfrak{g}_{\mathbb{R}}$. The restriction of $\kappa_{\mathbb{R}}(-,-)$ to $\mathfrak{h}_{\mathbb{R}}$ is positive definite. In particular, $\kappa_{\mathbb{R}}(-,-)$ identifies $\mathfrak{h}_{\mathbb{R}}$ with $\mathfrak{h}_{\mathbb{R}}^{*}$ and there is an induced inner product $\kappa^{*}(-,-)$ on $\mathfrak{h}_{\mathbb{R}}^{*}$. Write $\langle-,$-$\rangle_{\mathbb{R}}$ for the natural pairing between $\mathfrak{h}_{\mathbb{R}}^{*}$ and $\mathfrak{h}_{\mathbb{R}}$.
Let $\mathfrak{g},\mathfrak{k},\mathfrak{p}$, and $\mathfrak{h}$ denote the complexifications of $\mathfrak{g}_{\mathbb{R}}$, $\mathfrak{k}_{\mathbb{R}}$, $\mathfrak{p}_{\mathbb{R}}$, and $\mathfrak{h}_{\mathbb{R}}$. We have $\theta=\mathbb{C}\otimes\theta_{\mathbb{R}}$, an order two automorphism of $\mathfrak{g}$ such that $\mathfrak{g}^{\theta}=\mathfrak{k}$ and $\mathfrak{p}$ is the 1-1 eigenspace of $\theta$. We also write $\kappa(-,-)$ and $\langle-,$-$\rangle$ for the complexifications of $\kappa_{\mathbb{R}}(-,-)$ and $\langle-,$-$\rangle_{\mathbb{R}}$.
@@ 2.2. Roots and generators
For $\alpha\in R$, we denote by $t_{\alpha}\in\mathfrak{h}$ the element corresponding to $\alpha\in\mathfrak{h}^{*}$ under the identification $\mathfrak{h}\xrightarrow{\cong}\mathfrak{h}^{*}$ induced by $\kappa$, i.e., $t_{\alpha}\in\mathfrak{h}$ is the unique element such that
@@
This means $\alpha$ is real valued on $\mathfrak{h}_{\mathbb{R}}^{*}$ and $t_{\alpha}\in\mathfrak{h}_{\mathbb{R}}$. We fix generators $E_{\alpha}\in\mathfrak{g}_{\mathbb{R}}\cap\mathfrak{g}_{\alpha}$, for all $\alpha\in R$, such that $E_{\alpha},E_{-{\alpha}}]=t_{\alpha}$, i.e., $\kappa_{\mathbb{R}}(E_{\alpha},E_{-\alpha})=1$, and $\theta_{\mathbb{R}}(E_{\alpha})=-E_{-\alpha}$ (it is possible by [Humphreys, §25.2]). Note that
@@
$E_{\alpha},F_{\alpha}]=t_{\alpha}.$
@@ 2.3. Iwasawa decomposition
For $\mathfrak{sl}_{2}$, we denote the generator of $\mathfrak{k}=\mathfrak{so}_{2}$ simply by B:=FEB:=F-E.
@@ 2.4. Chevalley involution and generators
Thus, $\theta$ is conjugate to the standard Chevalley involution, which swaps $e_{\alpha_{i}}$ and $f_{\alpha_{i}}$ and acts as 1-1 on $\mathfrak{h}$, the existence of which is a Corollary of Serre’s presentation of $\mathfrak{g}$. In general, Chevalley involution refers to the automorphism of $\mathfrak{g}$ corresponding to the Cartan involution of the split real form.
@@ 2.5. Integrable representations
The Lie algebra $\mathfrak{k}$ is reductive which follows from the classification result [KnappLieGroups, Appendix C]. Fix a connected reductive group $K$ with the Lie algebra $\mathfrak{k}$. We call a $\mathfrak{k}$-representation integrable if it comes from a representation of $K$. By a “finite-dimensional representation of $\mathfrak{g}$” we understand a Harish-Chandra $(\mathfrak{g},K)$-module, i.e., a $\mathfrak{g}$-module which is integrable over $\mathfrak{k}$. In what follows, we will not refer to $K$ explicitly.
@@ 2.6. Weights and representations
This is a non-degenerate $\mathfrak{g}$-invariant form on $\mathfrak{g}$. We also write (,)(-,-) to denote the induced form on $\mathfrak{h}^{*}$, which is $W$-invariant and for any short root $\alpha$ satisfies
@@
In fact, since $W$ acts irreducibly on $\mathfrak{h}^{*}$ these two properties characterize (,)(-,-).
@@ 3.1. Braid group action for Weyl group
lying in a certain completion of $\mathrm{U}(\mathfrak{sl}_{2})$. Its action is well-defined on any finite-dimensional $\mathfrak{sl}_{2}$-module $V$. Moreover, the Weyl group operator is group-like, i.e., $\Delta(\tilde{s}^{\mathfrak{sl}_{2}})=\tilde{s}^{\mathfrak{sl}_{2}}\otimes\tilde{s}^{\mathfrak{sl}_{2}}$.
@@ 3.2. Braid group action for boundary Weyl group: $\mathfrak{so}_{2}\subset\mathfrak{sl}_{2}$
Consider the $\mathfrak{sl}_{2}$-module $V(1)$, with weight basis $v_{+}$ and $v_{-}$. We have $b=f-e\in\mathfrak{so}_{2}$ acting by $b\cdot v_{+}=v_{-}$ and $b\cdot v_{-}=-v_{+}$. Thus, $V(1)|_{\mathfrak{so}_{2}}=\mathbb{C}\{v_{+}-iv_{-}\}\oplus\mathbb{C}\{v_{+}+iv_{-}\}\cong\mathbb{C}_{i}\oplus\mathbb{C}_{-i}$. In this representation, we have $\tilde{s}=b$, so $\tilde{s}$ acts on $\mathbb{C}_{i}$ by $i$ and on $\mathbb{C}_{-i}$ by i-i.
@@
For any finite-dimensional module $V$ of $\mathfrak{sl}_{2}$, the action of $\tilde{s}^{\mathfrak{sl}_{2}}$ is equal to the action of $\exp\left(\frac{\pi b}{2}\right)$ understood in the sense of ˜3.4.
@@
By Example˜3.3, the actions agree on $V(1)$. Since $\Delta(b)=b\otimes 1+1\otimes b$ is a sum of two commuting operators, the operator $\exp\left(\frac{\pi b}{2}\right)$ is group-like, i.e., $\Delta(\exp\left(\frac{\pi b}{2}\right))=\exp\left(\Delta(\frac{\pi b}{2})\right)=\exp\left(\frac{\pi b}{2}\right)\otimes\exp\left(\frac{\pi b}{2}\right)$. Since the operator $\tilde{s}$ is group-like as well, their actions agree on any tensor product $V(1)^{\otimes n}$. As any irreducible representation of $\mathfrak{sl}_{2}$ can be realized as a subrepresentation of a suitable tensor product, the claim is then a consequence of complete reducibility of finite dimensional $\mathfrak{sl}_{2}$-modules. ∎
@@
[…]
@@
We see that $s_{0}(2x)$ and $is_{1}(2x+1)$ satisfy the same inductive relations as $\tilde{s}^{\mathfrak{so}_{2}}(2x)$ and $\tilde{s}^{\mathfrak{so}_{2}}(2x+1)$. Therefore, the statement follows from the equalities: $\tilde{s}^{\mathfrak{so}_{2}}(0)=s_{0}(0)=1$ and $\tilde{s}^{\mathfrak{so}_{2}}(1)=-is_{1}(1)=-i$. ∎
@@ 3.3. Braid group action for boundary Weyl group: $\mathfrak{k}\subset\mathfrak{g}$
Similarly to the usual Weyl group operators, we define the boundary Weyl group for the split symmetric pair $(\mathfrak{g},\mathfrak{k})$ in terms of simple reflection. In what follows, we are in the setting of Section˜5.
@@
Let $\alpha\in\mathfrak{h}^{*}$ be a simple root. Define the boundary Weyl group operator $\tilde{s}^{\mathfrak{k}}_{\alpha}:=\exp\left(\frac{\pi b_{\alpha}}{2}\right)$, understood in the sense of ˜3.4. When $\alpha=\alpha_{i}$, we will also denote $\tilde{s}^{\mathfrak{k}}_{i}:=\tilde{s}^{\mathfrak{k}}_{\alpha_{i}}$.
@@
For a finite dimensional $\mathfrak{g}$-module $V$, we have the usual Weyl group operators $\widetilde{w}^{\mathfrak{g}}$ acting on $V$. By Proposition˜3.5, the action of $\tilde{s}^{\mathfrak{k}}_{i_{1}}\ldots\tilde{s}^{\mathfrak{k}}_{i_{l}}$ on $V|_{\mathfrak{k}}$ coincides with the action of $\widetilde{w}^{\mathfrak{g}}$. In particular, if $X$ is an integrable $\mathfrak{k}$-module and $V$ is a finite dimensional $\mathfrak{g}$-module, then, recalling the conventions for $\Delta$ set forth in Section 2.5, we find
@@
Since the operators $\tilde{s}_{\alpha}^{\mathfrak{g}}$ satisfy the braid relations, it is enough to prove the proposition for integrable representations $X_{1},\ldots,X_{l}$ of $\mathfrak{k}$ such that any other irreducible representation can be obtained as a direct summand of $X_{i}\otimes V$ for some finite dimensional $\mathfrak{g}$-representation $V$. Moreover, it is enough to prove this for rank 2 algebras. This is done in Appendix˜B. ∎
@@
Moreover, Proposition˜3.5 implies $\mathrm{Ad}_{\tilde{s}^{\mathfrak{g}}_{i}}=\mathrm{Ad}_{\tilde{s}^{\mathfrak{k}}_{i}}$, hence $\mathrm{Ad}_{\tilde{s}^{\mathfrak{g}}_{i}}$ commutes with the Cartan involution $\theta:\mathfrak{g}\rightarrow\mathfrak{g}$. Therefore,
@@ 4.1. Fusion operator
Since $\mathfrak{h}$ is an abelian Lie algebra, we can identify its universal enveloping algebra $\mathrm{U}(\mathfrak{h})$ with the algebra $\mathcal{O}(\mathfrak{h}^{*})$ of polynomial functions on the dual space $\mathfrak{h}^{*}$. In what follows, we will often need to restrict to a certain open subset of $\mathfrak{h}^{*}$.
@@
We denote by $\mathcal{M}(\mathfrak{h}^{*})$ the space of meromorphic functions on $\mathfrak{h}^{*}$ with poles on the hyperplanes $\{\langle\alpha^{\vee},\lambda\rangle=k\}$ for all $\alpha\in R$ and $k\in\mathbb{Z}$. In particular, we have an algebra embedding $\mathrm{U}(\mathfrak{h})\cong\mathcal{O}(\mathfrak{h}^{*})\hookrightarrow\mathcal{M}(\mathfrak{h}^{*})$.
@@
Let $V_{1},V_{2},\dots,V_{d}$ be $\mathfrak{g}$-representations. For a function $f:\mathfrak{h}^{*}\rightarrow\mathrm{End}(V_{1}\otimes V_{2}\otimes\cdots\otimes V_{d})$, denote by $\lambda\mapsto f(\lambda-h|_{V_{i}})$, the function
@@
Let $V$ be a finite-dimensional representation of $\mathfrak{g}$. For a weight $\nu\in\mathfrak{h}^{*}$, we denote by $V[\nu]$ the corresponding weight space of $V$. Consider the space of intertwiners $\mathrm{Hom}_{\mathfrak{g}}(M_{\lambda},M_{\mu}\otimes V)$.
@@
    $\Phi^{v}_{\lambda}(m_{\lambda})=\Xi(\lambda)(m_{\mu}\otimes v).$
@@
The second assertion follows directly from [EtingofStyrkas, Proposition 2.1]: for every $\mu$, the singular vectors in $M_{\mu}\otimes V$ satisfying (4.3) can be obtained by applying the inverse of the Shapovalov form. Since the matrix coefficients of the latter are rational functions in $\mu$ and $\mu=\lambda-h|_{V}$ in the notations of ˜4.4, we conclude. ∎
@@
defines an intertwiner. Assuming the weights $\lambda$, $\mu$, and $\nu$ are generic, this intertwiner can be identified with $\Phi_{\lambda}^{u}$ for some $u\in(W\otimes V)[\lambda-\nu]$.